OpenStudy (anonymous):

what is the integral of [ t^9 sin(1-t^5) dt ]

OpenStudy (astrophysics):

this is super easy..for sat

OpenStudy (anonymous):

lol hell no

OpenStudy (astrophysics):

I'd use by parts

OpenStudy (anonymous):

i was thinking u - sub then parts, but perhaps best is to use the "subtraction" formula for \(\sin(1-t^5)\)

OpenStudy (astrophysics):

Yeah you would need a u sub at first though

OpenStudy (anonymous):

that would be messy also

OpenStudy (anonymous):

think it might be easier to write it as \[t^9\sin(1)\cos(t^5)-t^9\cos(1)\sin(t^5)\]

OpenStudy (anonymous):

if you start with \(u=1-t^5\) you have to solve for \(t^9\) in terms of \(u\) which is going to be a mess

OpenStudy (astrophysics):

I think it's messy either way

OpenStudy (astrophysics):

Actually no I think yourmethod is better

OpenStudy (anonymous):

no this way you have \(u=t^5\) so first one is \[\frac{\sin(1)}{5}\int u\cos(u)du\]

OpenStudy (anonymous):

i should shut up i hate this stuff

OpenStudy (astrophysics):

u=t^5 is a good sub

OpenStudy (astrophysics):

but it will still be messy xD

OpenStudy (astrophysics):

\[\int\limits \frac{ 1 }{ 5 } usin(1-u)du\]

OpenStudy (astrophysics):

Haha I think you would do another u sub then by parts omg

OpenStudy (anonymous):

yeah that is why it is probably easiest to use subtraction angle formula first

OpenStudy (astrophysics):

Yeah haha, ok op you have the techniques, good luck!

OpenStudy (anonymous):

\[\sin(1)\int t^9\cos(t^5)dt-\cos(1)\int t^9\sin(t^5)dt\]is where i would start

zepdrix (zepdrix):

U-sub into parts doesn't seem that bad :p Only requires a single parts application because of how t^9 breaks down. \(\large\rm u=1-t^5\qquad\qquad\to\qquad\qquad 1-u=t^5\) \(\large\rm -\frac{1}{5}du=t^4dt\) And use those three pieces to give you\[\large\rm -\frac15\int\limits(1-u)\sin u~du\] And then parts, bam bam bam :D

zepdrix (zepdrix):

@dramirez9 what do you think? :) Figure this one out yet?

OpenStudy (anonymous):

@zepdrix THANKS!!!!!! it was becoming a nightmare using substitution by parts.

zepdrix (zepdrix):

Got it? cool cool

OpenStudy (astrophysics):

\c:/